### v1

# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

Abstract
The European electricity market is based on large pricing zones with a uniform day-ahead price. The energy transition leads to shifts in supply and demand and increasing redispatch costs. In an attempt to ensure efficient market clearing and congestion management, the EU Commission has mandated the Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. Based on a unique data set published in the context of the BZR, we compare various pricing rules for the German power market. We compare market clearing and pricing for national, zonal, and nodal models, including their generation costs and associated redispatch costs. Moreover, we investigate different non-uniform pricing rules and their economic implications for the German electricity market. Our results indicate that the differences in the average prices in different zones are small. The total costs across different configurations are similar and the reduction of standard deviations in prices is also small based on this data set. A nodal pricing rule leads to the lowest total costs. We also analyze the quality of different pricing rules and their differences with respect to the quality of the price signals and the necessary uplift payments. While the study focuses on Germany, the analysis is relevant beyond and feeds into the broader discussion about pricing rules.
[…]
[1]organization=School of Computation, Information and Technology, Technical University of Munich,addressline=Boltzmannstr. 3, city=Garching, postcode=85748, country=Germany
1 Introduction
A fundamental problem in electricity market design is to compute adequate price signals in the presence of non-convex preferences of market participants. Prices should reflect the true value of electricity and provide the right bidding and investment incentives to all relevant market participants (Cramton,, 2017). In liberalized markets, day-ahead electricity prices are derived from welfare maximization problems, but pricing is challenging due to the non-convexity of these optimization problems (Liberopoulos & Andrianesis,, 2016). A central discussion on electricity market pricing revolves around zonal and nodal pricing (aka. locational marginal pricing). Among the electricity wholesale markets implementing nodal pricing are Argentina, Chile, Mexico, New Zealand, Peru, Russia, Singapore, and several regions in the United States (U.S.). In contrast, a zonal pricing approach has been employed in the large and coupled European markets to reduce the computational burden of market clearing and pricing. Large groups of nodes are aggregated into zones, and the market clearing problem only considers flow constraints between zones. With some exceptions (e.g., Italy, Norway, Sweden), there is a single price zone within a country, and the day-ahead market provides a uniform national electricity price. Ignoring transmission constraints can lead to dispatch decisions that are not physically feasible. Thus, after the market clearing, transmission operators conduct redispatch to ensure that the final allocation aligns with physically feasible power flows.
1.1 Zonal vs. Nodal Pricing
Due to climate change and the ongoing energy transition, the zonal pricing paradigm has come under scrutiny (Eicke & Schittekatte,, 2022; Bertsch et al.,, 2016; Trepper et al.,, 2015). Volatile renewable energy sources such as wind and solar lead to congestion in different parts of the electricity grid (Neuhoff et al.,, 2013). In addition, the energy transition can lead to structural problems. We will focus on Germany as a case in point. Wind farms have grown significantly in Northern Germany, while much of the industrial demand is in the south. Energy is cleared as if there were no transmission constraints, but the traded energy can often not be delivered. As a result, redispatch costs have increased drastically in recent years to 4.2 billion in 2022 (Bundesnetzagentur,, 2022). The redispatch costs in 2022 were also influenced by the war in the Ukraine, but have been increasing consistently over several years. In an attempt to ensure efficient market clearing and congestion management, the EU Commission – as part of the Clean Energy Package – has mandated a Bidding Zone Review (BZR) process to reevaluate the configuration of European bidding zones. An integral part of this process is a locational marginal pricing (LMP) study conducted by the European Network of Transmission System Operators for Electricity (ENTSO-E) as a basis to identify structural congestion and modified bidding zones. Based on this study, in August 2022, the European Union Agency for the Cooperation of Energy Regulators (ACER) decided on a set of alternative bidding zone configurations based on proposals by ENTSO-E (ACER, 2022a, ). Among others, the BZR considers a split of the German price zone.
In their LMP study, ENTSO-E solved linearized unit commitment models with marginal pricing to obtain nodal prices (ENTSO-E,, 2022). The data is the main input for the Bidding Zone Review and constitutes the results of a multi-year effort of the European TSOs. The discussion of nodal and zonal prices has a long history, but this unique data set allows for a unique opportunity to compare different pricing rules on realistic data and estimate the differences. Note that the allocation considers assets individually at each node and abstracts from block orders or portfolio bids that are common practice in Europe today. The computed prices served as input for clustering approaches to yield alternative bidding zone configurations. Following Article 16 of the BZR methodology, ENTSO-E has published non-confidential data related to the LMP study (ENTSO-E,, 2023). We leverage this dataset to expand on the results of the LMP study, focusing on the German bidding zone. Beyond the choice between four alternative zonal configurations of the German market, our goal is to understand how these configurations compare to nodal prices and the current market with a uniform national price.
Specifically, we compare market clearing and pricing for national, zonal, and nodal models, the generation costs and associated redispatch costs. We consider a redispatch model that minimizes cost-based compensations provided to generators for any changes in their dispatched quantities during redispatch. Moreover, we investigate different non-uniform pricing rules and their economic implications for the German electricity market. We consider Integer Programming (IP) pricing (O’Neill et al.,, 2005) and Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) as two established pricing rules in U.S. markets, and the recently suggested Join pricing rule (Ahunbay et al.,, 2022). We also implemented a version of the Euphemia algorithm currently used in European day-ahead markets. In contrast to IP, CH, and Join pricing, Euphemia sacrifices welfare maximization to achieve linear and anonymous prices and budget balance. We compare the outcomes of clearing and pricing rules based on the total costs, average prices, make-whole payments (to compensate paradoxically accepted bids), and lost opportunity costs. For Euphemia, we also compute the welfare loss incurred by the algorithm. We report aggregate results using the generation and demand scenarios for 2025 and climate year data of 2009 provided by the ENTSO-E LMP study. We also focus on individual days to illustrate our findings and their economic implications.
1.2 Contributions
Our study suggests that a nodal pricing rule lowers system costs as costly redispatch can be avoided. Conversely, by disregarding transmission constraints, the zonal configurations lead to an inefficient dispatch. Even if redispatch is conducted at minimal costs, this inefficiency leads to higher overall system costs. The average price differences across configurations between 1 and 4 zones for Germany are less than 3 EUR/MWh. This could partly be attributed to the capacities assumed by ENTSO-E for 2025, e.g., new solar capacities in the southern region. There are also individual days where the price difference between zones is more than 10 EUR/MWh, but on average the differences are small. The higher total costs for national and zonal prices arise from the fact that due to the omission of network constraints, an inefficient allocation is picked. This results in the selection of generators that would not be part of an efficient nodal dispatch. These inefficiencies lead to higher total costs after redispatch that is required to achieve a feasible dispatch respecting the transmission capacities. As indicated earlier, in the past there was often an oversupply of renewable energy sources in the north and high demand in the south of Germany, which raised expectations that redispatch decreases significantly with two price zones. However, our results based on data from ENTSO-E, does not confirm such expectations.
The redispatch costs in our model do not decrease considerably with a split in 2, 3, or 4 zones. The way how these redispatch costs are computed can be considered as a lower bound. For example, we assume that every generator can be redispatched, but this is not the case in reality. A key goal for the Bidding Zone Review is to find pricing zones such that the variance in prices is small, which is seen as a proxy for network congestion. We show that the differences in the standard deviation of prices caused by congestion in different zones are small compared to the overall standard deviation in prices, which explains why the differences in the redispatch costs are small across the different zonal configurations.
We also provide an analysis of different non-uniform pricing rules such as IP pricing (as used in U.S. power markets), Convex Hull Pricing, and the Join pricing rule. The results illustrate the trade-offs that can be expected in terms of quality of the congestion signals and minimum make-whole payments and provide guidance regarding the analysis of non-uniform pricing rules in the context of CACM 2.0 (All NEMO Committee, 2023b, ). While our analysis focuses on Germany, the policy implications go beyond. The study highlights the challenges in delineating price zones in a zonal market and the trade-offs between different pricing rules based on a new and unique data set that results from a multi-year effort by the European TSOs.
1.3 Limitations and Organization
Like any empirical analysis of its kind, also our study faces limitations. First, our examination focuses on the German day-ahead market without incorporating neighboring countries, cross-border trades, or loop flows through other countries. Second, we do not model the intraday or forward markets, which hold considerable significance within the European electricity system. Third, we assume fixed demand from the ENTSO-E dataset, neglecting the potential for demand reduction incentives under nodal prices. Moreover, the current European market is characterized by portfolio bidding, differing from the unit-commitment bids underlying both the study by ENTSO-E and our analysis. Portfolio bids allow optimization of trades across assets and differ from those feasible in nodal market designs. Incorporating such considerations into our analysis would require making strong and potentially unwarranted assumptions regarding market participant behavior. Finally, the data set provided by ENTSO-E was constructed for the target year 2025, and even though different climate years are considered, it is important to acknowledge that analyses can vary based on differing assumptions regarding supply and demand. However, the ENTSO-E dataset represents the best source available today for such a study. Additional assumptions made for our analysis are described in Section 3. We also emphasize that the discussion surrounding zonal and nodal pricing extends beyond dispatch and prices. In the ongoing Bidding Zone Review, a comprehensive approach is adopted which considers factors such as liquidity, security of supply, transition costs, and investments in low-carbon technologies, which we do not in our analysis.
The remainder of this article is structured as follows. In Section 2, we introduce our unit commitment model and formalize nodal and zonal market clearing. We also review the IP, CH, and Join pricing rules. Section 3 elaborates on the data sources used to implement our dispatch model and outlines our experimental design. Section 4 presents the key results of our analysis of different market clearing and pricing rules. Lastly, Section 5 provides a summary and conclusions.
2 Clearing and Pricing on Electricity Markets
2.1 Preliminaries
We consider an electricity market consisting of a set of buyers $B$ and a set of sellers $S$, each located at nodes $N$ in an interconnected electricity network. The set of power lines $L$ is encoded as pairs of nodes $(n,m)$, and we consider multiple periods $T$. An item in this market corresponds to a unit of electricity at a particular node $n\in N$ at a specific time $t\in T$.
A buyer $b\in B$ possesses a valuation function $v_{b}:\mathbb{R}^{N\times T}\rightarrow\mathbb{R}$ that quantifies the buyer’s preferences and constraints. Similarly, a seller $s\in S$ has a cost function $c_{s}:\mathbb{R}^{N\times T}\rightarrow\mathbb{R}$.
The market operator collects these buy and sell bids to calculate a feasible allocation. As electricity is transmitted over the network, the allocation is subject to physical power flows, encoded as a constraint set $\Psi$. An accurate representation of the transmission network would require $\Psi$ to be equivalent to AC optimal power flow (ACOPF) constraints (Molzahn & Hiskens,, 2019). However, it is well known that the ACOPF is intractable for realistic problem sizes, and electricity market operators need to employ computationally scalable approximations of the ACOPF.
2.2 Market Clearing
In practice, market operators employ different bidding languages to encode $v_{b}$ and $c_{s}$, as well as different approximations $\Psi$ of the power flow constraints. In this work, we use the data released for the European bidding zone review (BRZ) to compare allocation and prices under different market clearing mechanisms.
In full generality, the market operator seeks an allocation $(x,y)$, where $x=(x_{b})_{b\in B}$, $x_{b}\in\mathbb{R}^{N\times T}$ is the allocation vector of buyers and $y=(y_{s})_{s\in S}$,$y_{s}\in\mathbb{R}^{N\times T}$ is the allocation vector of sellers. The market operator seeks an allocation that maximizes welfare, by solving the following optimization model:
$\max_{x,y}\quad$ $\sum_{b\in B}v_{b}(x_{b})-\sum_{s\in S}c_{s}(y_{s})$ (1)
subject to $x,y\in\Psi.$
On European day-ahead markets, buyers and sellers submit hourly bids and block orders (NEMO Committee,, 2019). The BZR data, however, does not contain bids of this structure. Instead, it provides generator and load characteristics similar to unit commitment problems in U.S. markets. Specifically, for buyers $b$ there is only information on a fixed demand profile $P_{b}\in\mathbb{R}^{N\times T}$, while price-elastic bids are unavailable. Consequently, the valuation function of $b$ simplifies to strict demand satisfaction.
$v_{b}(x_{b})=\begin{cases}-\infty,\;\;x_{b}\neq P_{b}\\ 0,\;\;\;x_{b}=P_{b}\end{cases}$ (2)
For sellers/generators $s\in S$, the BZR data release includes more detailed information. Each generator $s$ has a minimum and maximum production quantity $\underline{P}_{s}\in\mathbb{R}$ and $\overline{P}_{s}\in\mathbb{R}$ at their node $n_{s}$. Moreover, once a generator has been started, there is a minimum uptime constraint to run for at least $\underline{R}_{s}\in\mathbb{Z}_{0}^{+}$ periods. Generators incur variable costs $g_{s}$ to produce electricity and fixed costs $h_{s}$ whenever the generator runs. We use a binary commitment variable $u_{st}\in\{0,1\}$ to model a generator’s constraint set $P_{s}$.
$P_{s}=\{y_{s}\in\mathbb{R}^{N\times T}:\;$ $y_{s,n_{s}t}\geq\underline{P}_{st}u_{st}$ $\;\forall t\in T,$
$y_{s,n_{s}t}\leq\overline{P}_{st}u_{st}$ $\;\forall t\in T,$
$\phi_{s,n_{s}t}\geq u_{st}-u_{s(t-1)}$ $\;\forall t\in T\setminus\{T_{0}\},$
$\sum_{i=t-\underline{R}_{s}+1}^{t}\phi_{s,n_{s}i}\leq u_{st}$ $\;\forall t\in T\setminus\{T_{0}\},$
$y_{s,nt}=0$ $\;\forall n\in N\setminus\{n_{s}\},t\in T$
$u_{st}\in\{0,1\}$ $\;\forall s\in S,t\in T.\}$
With this constraint set, we define a generator’s cost function $c_{s}$ as follows.
$c_{s}(y_{s})=\begin{cases}\infty,\;\;y_{s}\notin P_{s}\\ \sum_{t\in T}g_{s}y_{s,n_{s}t}+\sum_{t\in T}h_{s}u_{st},\;\;y_{s}\in P_{s}\end{cases}$ (3)
During the 2000s and 2010s, many U.S. ISO markets moved from zonal to nodal pricing. Recognizing that an accurate representation of the transmission network in the form of the ACOPF is computationally infeasible, market operators resort to linearized versions of power flow. The most common implementation is the Direct Current OPF (DCOPF), which is based on three simplifying assumptions (Stott et al.,, 2009; Molzahn & Hiskens,, 2019), i.e., (1) uniform voltage magnitudes, (2) negligibly small voltage angle differences, and (3) neglecting line resistances and reactive power. Under these assumptions, the flow constraints can be expressed as simple linear constraints.
$\Psi^{DC}=\{(x,y):\sum_{s\in S}y_{s,nt}-\sum_{b\in B}x_{b,nt}=\sum_{m\in N}-B_{nm}(\theta_{n}-\theta_{m})\;\forall n\in N,t\in T\}$ (4)
In (4), $B_{nm}$ is the susceptance of the line connecting nodes $n$ and $m$. If such a line does not exist, then $B_{nm}=0$, and $B_{nn}=0\;\forall n$. The voltage angle decision variable for node $n$ is denoted as $\theta_{n}$. The constraint set in (4) can be enriched by further constraints, such as thermal line limits, box constraints on the voltage angles, and more (Li et al.,, 2022). We use such a standard DCOPF model in our analysis, similar to what was used in the ENTSO-E study. When the representation $\Psi$ of the transmission network considers all nodes in the transmission network, we refer to nodal market clearing. When only a subset of the grid is considered, we speak of zonal market clearing, as discussed in the following.
2.2.1 Zonal Market Clearing
The key idea of zonal pricing is to aggregate nodes into zones and only consider transmission flows across zones. This simplifies the constraint set $\Psi$ significantly, enhances computational scalability, and allows for portfolio bidding of larger market participants. In the European market, two approaches have been implemented: the available transfer capacity (ATC) model and the flow-based market coupling (FBMC) methodology (NEMO Committee,, 2019). Both approaches require a refined capacity calculation process before the allocation to assess the available cross-zonal flow capacities. Furthermore, the FBMC approach necessitates access to a power transfer distribution factor (PTDF) matrix that indicates how zone net position changes affect cross-zonal flows.
As these data were not included in the BZR data release, we rely on a different approach to compute a zonal allocation. Specifically, to determine the aggregate characteristics of a cross-zonal connection, we sum up the thermal limits and use the mean susceptance over all relevant cross-zonal lines. Specifically, with a set of zones $Z$ and $B_{zz^{\prime}}=\overline{B}_{nm:n\in z,m\in z^{\prime}}$ as the mean cross-zonal susceptance, the zonal constraint set is defined as follows.
$\Psi^{DC}_{Zonal}=\{(x,y):\sum_{s\in S}y_{s,zt}-\sum_{b\in B}x_{b,zt}=\sum_{z^{\prime}\in Z}-B_{zz^{\prime}}(\phi_{z}-\phi_{z}^{\prime})\;\forall z\in Z,t\in T\}$ (5)
In essence, (5) is a natural extension of (4) where a set of aggregate zones is considered instead of nodes. Cross-zonal flows are feasible if they satisfy $\Psi^{DC}_{Zonal}$.
The mapping of nodes to zones is suggested by the ENTSO-E study on locational marginal pricing (ACER, 2022b, ). Four different zonal splits were proposed for the German bidding zone, suggesting between two to four price zones.
2.2.2 Redispatch
While zonal market clearing is computationally less expensive than nodal market clearing, it substantially simplifies the underlying physical power flows. In fact, the economic outcome of zonal market clearing is usually infeasible with physical power flows. As a result, transmission operators resort to redispatch, i.e., modifying the economic allocation to provide a physically feasible outcome.
Current redispatch mechanisms consider two paradigms: Firstly, cost-based redispatch, as implemented in Germany, reimburses redispatched generators for their additional costs or lost profits and makes them indifferent to the previous market outcome. Secondly, market-based redispatch aims at a bid-based procurement of redispatch volumes. Market-based redispatch was criticized as prone to market power abuse and causing inc-dec gaming, i.e., strategic overbidding in day-ahead markets to generate higher profits in subsequent redispatch markets (Hirth & Schlecht,, 2020).
This work simulates a cost-based redispatch in line with real-world practices. As redispatch should be cost-minimal, the fundamental idea of our approach is to find a modified allocation that satisfies the constraint set of $\Psi^{DC}$ and, at the same time, minimizes the required compensations to sellers for their changes in generation. Demand is excluded from redispatch, as it is current practice in Germany. More formally, with $(x^{Zonal},y^{Zonal})$ as the outcome of the zonal allocation, we solve the following problem:
$\min_{x,y}\quad$ $\sum_{s\in S}\lvert c_{s}(y_{s})-c_{s}(y_{s}^{Zonal})\rvert$ (6)
subject to $x=x^{Zonal}$
$x,y\in\Psi^{DC}$
The objective of the optimization problem described in (6) is to adjust the zonal allocation $(x^{Zonal},y^{Zonal})$ in a manner that yields a DCOPF-feasible solution $(x,y)$ at minimal costs. This adjustment involves compensating sellers for any changes in their dispatched quantities during redispatch. For each seller $s$, this compensation is determined by the difference between the costs they incur under the initial zonal allocation $c_{s}(y_{s}^{Zonal})$ and the costs under the final DCOPF-feasible allocation $c_{s}(y_{s})$. The objective of (6) is to minimize the total costs associated with these adjustments. As part of our experiments we also report results for redispatch that minimizes volumes instead of costs, i.e., the redispatch objective function changes to $\min_{x,y}\;\sum_{s\in S}\lvert y_{s}-y_{s}^{Zonal}\rvert$.
Problem (6) yields the final DCOPF-feasible allocation $(x,y)$ that can be implemented as physical power flows. Note that a DCOPF-feasible solution may still be ACOPF-infeasible and require further adjustments to the allocation, which we disregard in this work. Market participants subject to redispatch must be compensated for lost profits, reducing overall welfare. In our numerical experiments, we thus consider system costs as the total cost of generation before redispatch, redispatch costs themselves, and total costs as the sum of system and redispatch costs.
2.3 Pricing on Non-Convex Markets
Once an allocation $(x,y)$ has been obtained, the market operator must provide electricity prices $p\in\mathbb{R}^{N\times T}$ for each node and time period. Due to the non-convexities implied by sellers’ cost functions (i.e., the binary commitment variables $u_{st}$), this is not a trivial task. In microeconomics, the welfare theorems provide a foundation for pricing in markets. In their seminal paper, Arrow & Debreu, (1954) demonstrate that under convex preferences, demand independence, and perfect competition with divisible items, a market operator can always achieve a set of Walrasian equilibrium prices that support the welfare-maximizing allocation. We assume quasilinear utilities, i.e., the utility of each market participant is defined as the difference between valuation or costs and price.
$u_{b}(x|p)=v_{b}(x)-\sum_{n\in N,t\in T}p_{nt}x_{b,nt}$ $\quad\forall\ b\in B$ (7)
$u_{s}(y|p)=\sum_{n\in N,t\in T}p_{nt}y_{s,nt}-c_{s}(y)$ $\quad\forall\ s\in S$
At given prices $p$, each market participant has some preferred bundles that maximize utility. We call these bundles the demand set and denote the maximum possible utility as follows.
$\hat{u}_{b}(p)=\max_{x}v_{b}(x)-\sum_{n\in N,t\in T}p_{nt}x_{b,nt}$ $\quad\forall\ b\in B$ (8)
$\hat{u}_{s}(p)=\max_{y}\sum_{n\in N,t\in T}p_{nt}y_{s,nt}-c_{s}(y)$ $\quad\forall\ s\in S.$
A Walrasian equilibrium consists of a market-clearing allocation and linear prices such that no participant has an incentive to deviate.
Definition 1 (Walrasian Equilibrium).
A price vector $p$ and a feasible allocation $(x,y)$ form a Walrasian equilibrium if:
  1. 1.
    (Market clearing) Supply equals demand, i.e., $\sum_{s\in S,n\in N}y_{s,nt}=\sum_{b\in B,n\in N}x_{b,nt}\;\forall t\in T$.
  2. 2.
    (Envy-freeness) Every market participant maximizes their utility at the prices, i.e., $u_{b}(x|p)=\hat{u}_{b}(p)\;\forall b\in B$ and $u_{s}(y|p)=\hat{u}_{s}(p)\;\forall s\in S$.
  3. 3.
    (Budget balance) The payments that sellers receive equal the payments that buyers provide, i.e., $\sum_{n\in N,t\in T}p_{nt}(\sum_{s\in S}y_{s,nt}-\sum_{b\in B}x_{b,nt})=0$.
Unfortunately, real-world electricity markets do not satisfy the assumptions underlying the welfare theorems. As in many other markets, items are not perfectly divisible and market participants exhibit non-convex preferences. A large stream of literature has investigated existence conditions for Walrasian equilibria in non-convex markets (Kelso & Crawford,, 1982; Bikhchandani & Ostroy,, 2002; Baldwin & Klemperer,, 2019), but none of these conditions are satisfied in electricity markets. Bikhchandani & Mamer, (1997) demonstrate that Walrasian equilibria exist if and only if the linear relaxation of the allocation problem has an integer solution. When this is not the case, market participants bear (global) lost opportunity costs at any set of prices and do not maximize their individual profits, violating envy-freeness.
Consequently, market operators have to resort to heuristic pricing rules that aim at approximating Walrasian equilibrium prices. Many such pricing rules have been proposed (Liberopoulos & Andrianesis,, 2016), each compromising on the properties of Walrasian equilibria differently. In this work, we focus on Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) and Integer Programming (IP) pricing (O’Neill et al.,, 2005) as two established pricing rules in practice, as well as on the Join pricing rule (Ahunbay et al.,, 2022) as a novel multi-objective approach to pricing. We also consider the Euphemia pricing rule as it is applied in European electricity markets.
2.3.1 Convex Hull Pricing
In the absence of Walrasian equilibrium prices, there is no set of prices $p$ that maximizes the profit of every market participant. The concept of global lost opportunity costs describes this forgone payoff.
Definition 2 (Global Lost Opportunity Costs (GLOCs)).
Given an allocation $(x,y)$ and prices $p$, a market participant’s global lost opportunity costs describe the difference between their individual payoff maximum, given $p$, and their actual payoff.
[…]
The key idea of CH pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) is to minimize GLOCs over all market participants. Formally, CH pricing replaces each non-convex cost function $c_{s}(y_{s})$ with its convex envelope in (1), and derives prices from the dual of the resulting convex problem.
Computing convex envelopes and determining CH prices is generally computationally challenging (Schiro et al.,, 2016). However, it has been shown by Hua & Baldick, (2017) that CH pricing becomes tractable when the cost function is described as in (3). Specifically, by relaxing the binary constraints $u_{st}\in\{0,1\}$ to $u_{st}\in[0,1]$ and solving the dual of (1), CH prices can be obtained from a single linear program. This approach, also referred to as Extended Locational Marginal Pricing (ELMP), was also used by the recent ENTSO-E study (ENTSO-E,, 2022).
2.3.2 IP Pricing
In a convex market, Walrasian equilibrium prices are equivalent to marginal prices, i.e., the costs of the last accepted bid set the prices. Following this notion, O’Neill et al., (2005) try to generalize marginal pricing for non-convex markets. Their IP pricing rule assumes that generators cannot easily deviate from their commitment status (as encoded by the binary variable $u_{st}$) and sets prices equal to the variable costs of the marginal committed unit (i.e., the marginal unit among all units with $u_{st}=1$).
IP pricing has gained popularity in many practical markets. It involves three steps: (i) obtaining the optimal commitment variables $u_{st}^{*}$ from the allocation problem in (1), (ii) fixing the commitment variables of each generator to these optimal values, i.e., setting $u_{st}\in\{0,1\}$ to $u_{st}\in[0,1]$ with $u_{st}=u_{st}^{*}$ in (3), and (iii) solving (1) with these linearized cost functions and extracting prices from its dual.
IP prices can be considered Walrasian equilibrium prices assuming that no generator can deviate from its commitment status. In other words, every participant maximizes their utility locally, meaning under fixed commitment. We define these local lost opportunity costs a subset of GLOCs.
Definition 3 (Local Lost Opportunity Costs (LLOCs)).
Given an allocation $(x,y)$, generator commitments $u^{*}$, and prices $p$, a seller’s local lost opportunity costs describe the difference between their individual payoff maximum under fixed commitment, given $p$, and their actual payoff.
[…]
By convexity of $v_{b}(x_{b})$, a buyer’s LLOCs exactly equal their GLOCs. Besides computational tractability, IP prices accurately reflect the marginal value of transmission capacity (Yang et al.,, 2019). Specifically, price differences among nodes arise solely when the network experiences congestion. It has been shown that this property immediately corresponds minimizing LLOCs (Ahunbay et al.,, 2022), making them a significant indicator of good congestion signals.
2.3.3 Join Pricing
In most electricity markets, neither GLOCs nor LLOCs are paid out to market participants to disincentivize deviations from the optimal allocation. Instead, market operators merely ensure that no participant incurs losses by participating in the market, ensuring non-negative utilities and, equivalently, individual rationality. The payments to ensure individual rationality are known as make-whole payments.
Definition 4 (Make-Whole Payments (MWPs)).
Given an allocation $(x,y)$ and prices $p$, a market participant’s make-whole payments describe their negative payoff, if applicable.
[…]
Bids that require MWPs are also known as paradoxically accepted bids. MWPs can be regarded as another subset of GLOCs, referring only to participants’ incentives to not participate in the market at all. However, both CH and IP pricing can experience high levels of MWPs. Recent practical concerns have thus motivated the development of pricing rules that ensure lower levels of MWPs. Among them is the Join pricing rule (Ahunbay et al.,, 2022), which proposes a dual pricing problem that combines the objective of IP pricing (minimizing LLOCs) with minimizing MWPs. Specifically, for each participant it considers the maximum of LLOCs and MWPs in the objective, leading to a robust and computationally tractable pricing rule with adequate congestion signals and low MWPs simultaneously.
2.4 Euphemia
CH, IP, and Join pricing all have in common that they price the welfare-maximizing allocation. Market participants incur lost opportunity costs; some are even paradoxically accepted and need to be compensated by MWPs.
European electricity markets follow a different approach. Paradoxically accepted bids are not permitted, and instead, the market operator deviates from the welfare-maximizing outcome to avoid paying any MWPs. This is the essence of the Euphemia day-ahead clearing algorithm (NEMO Committee,, 2019). Note that avoiding paradoxically accepted bids does not satisfy the envy-freeness property from Definition 1. Market participants still incur GLOCs and LLOCs, a subset of which are even rejected to participate in the market even though it would be profitable for them to do so. Such bids are referred to as paradoxically rejected bids.
The Euphemia algorithm combines market clearing and pricing (unlike the pricing rules above which can be applied to any pre-computed allocation). In European markets, it computes an allocation and zonal prices, but theoretically, it can also be applied in nodal markets. Algorithmically, it starts by computing the welfare-maximizing allocation, solving (1). It then advances to examining whether prices can be established for the allocation, i.e., they must prevent any paradoxically accepted bids, ensure the acceptance of all in-the-money hourly bids and, optionally, certify that cross-zonal flows occur from a lower-price to a higher-price zone. If paradoxically accepted bids exist, Euphemia adds cuts to (1) that eliminate the current candidate solution, constraining the allocation problem and leading to welfare losses. These iterations between allocation and pricing continue until a solution without paradoxically accepted bids is reached.
Subsequently, Euphemia continues by addressing some of the idiosyncrasies of the European market (PUN search, re-insertion of certain paradoxically rejected bids, resolving indeterminacy). If, at any of these steps, violations are detected, the iterative procedure continues. Eventually, Euphemia outputs allocation, zonal prices, and cross-zonal flows.
Since we aim to investigate prices and welfare losses associated with Euphemia, we implemented a simplified version of the algorithm for our numerical experiments. With the data provided, however, we can only replicate the first two steps of the algorithm, i.e., the iterative allocation and pricing procedure. When a paradoxically accepted seller $s$ in period $t$ is detected, we add a constraint $u_{st}=0$ to the allocation problem (1). No open-source implementation of the algorithm is available, but we followed the public description as closely as possible.
A practical problem with Euphemia is its computational tractability. Due to its iterative nature, multiple mixed-integer problems must be solved until allocation and prices are obtained. With the planned 15-min market time unit introduction in Europe, the scalability of Euphemia poses a significant concern, and policymakers consider non-uniform pricing rules (such as the ones introduced above) to obtain solutions faster (All NEMO Committee,, 2022).
3 Data and Processing
This study examines welfare, prices, GLOCs, LLOCs, and MWPs for multiple zonal/nodal configurations, allocation rules, and pricing mechanisms. As discussed, we leverage the data released in the context of the European bidding zone review (ENTSO-E,, 2022; ACER, 2022b, ). We focus on Germany as arguably one of the most important bidding zones regarding clearing volume and its central position in the European grid. In this section, we provide an overview of the data set and our processing of it.
3.1 Bidding Zone Review Data
In September 2022, ENTSO-E, on behalf of all European TSOs, released various data sets related to their locational pricing study (ENTSO-E,, 2022, 2023). The input data structure along two primary datasets: the grid model as a basis for the transmission network and the input files to model generation and demand. Data on reserves, storage, and imports/exports were not considered in our study. Along with these input data, results files such as locational prices and cleared generation, demand, and storage were released. The publication refers to a total of 24 weeks over three representative climate years (1989, 1995, and 2009). Only the climate information from 1989, 1995, and 2009 is used, while generation and demand scenarios were generated for the target year 2025 (ENTSO-E,, 2022).
Unfortunately, some parts of the data were aggregated before their publication (e.g., local demand to country-wide demand), and some data were not revealed (e.g., a mapping of generators in the input files to their location in the grid model). We elaborate on these issues in the following. Without access to these proprietary data, the outcome of the ENTSO-E study cannot be exactly reproduced. However, it still provides a foundation against which we can validate our own results.
The grid model for the German transmission system includes information on the system topology, components, and operational states. Each substation comprises multiple voltage levels, each of which can comprise several topological nodes that are connected with grid components / transmission lines. We consider power transformers as regular transmission lines and neglect their phase-shifting capability. For the grid constraints of our DCOPF in (4), we derive the per-unit susceptance $B_{nm}$ for each line from the given actual susceptance and base admittance (Grainger & Stevenson,, 1994). We also induce transmission limits on each power line, which we infer as the minimum of each line’s angular stability limit (inferred from its surge impedance loading and the St. Clair curve (St. Clair,, 1953)), voltage drop limit (inferred from voltage and reactance (Hao & Xu,, 2008)), and thermal limit (inferred from voltage and current on a three-phase system (Moreira et al.,, 2006)). We assume these transmission limits to be constant over time. Missing data, e.g., on the length and susceptance of lines, were replaced by the data in the JAO Static Grid Model where possible, and else inferred from the mean of all other lines.
The grid model consists of 2998 topological nodes. A significant part of these nodes represent breakers and disconnectors, and thus control devices that are neither associated with supply or demand nor changing voltage levels. Therefore, we disregard breakers and disconnectors and merge all topological nodes and their associated equipment at a given voltage level. Nodes not connected to the rest of the grid are merged with nodes associated with the same substation. After preprocessing, our grid model consists of 1687 nodes and 3232 lines (Schmitt,, 2023). In order to illustrate our results, we manually derived geographical coordinates for each substation from OpenStreetMap.
For the demand side, the input files provide hourly aggregated load profiles for the entire country of Germany. Neither the nodal distribution of demand nor demand valuations are provided. Demand-side response was considered in the ENTSO-E study, but no related data was released. Consequently, we distribute demand proportional to each consumer’s base load that is available as part of the grid model. We thereby assume that demand fluctuations over time occur uniformly across all consumers (Schmitt,, 2023). We further assume that demand is exclusively price-inelastic.
The input files pertaining to the supply side are more granular and separated by generator type, i.e., Hard Coal, Lignite, Gas, Light Oil, Solar, Hydro Run-Of-River, Hydro Reservoir, Hydro PumpOL, Hydro PumpCL, Onshore Wind, Offshore Wind, Other Non-RES, and Other RES. They contain operational characteristics of thermal units, derived from the Pan-European Market Modeling Database (PEMMDB), and hourly aggregated output of renewable energy sources, derived from Pan European Climate Database (PECD). The operational characteristics include – among others – minimum / maximum power and minimum runtime requirements that we use to parameterize our generator cost functions. The data also includes must-run obligations, start-up / shut-down times, and ramp rates, which may be used to further extend the model. Similarly, data on storage or demand response could leverage more detailed models.
As the maximum power of each seller is given as a fixed quantity (denoting its nominal capacity), we need to manually incorporate the variability of renewable energy sources. We thus distribute the aggregate dispatched energy in each hour equally to all plants of the same type, proportional to their nominal capacities, and set the maximum power to this value. The underlying assumptions are that renewables are always dispatched at maximum power and that meteorological conditions are equal across Germany. Because hydro power plants are dispatchable, they are excluded from this logic. The data also do not include information on variable or fixed costs for renewables, which we derived from literature (Kost et al.,, 2021; IRENA,, 2022; ENTSO-E,, 2022). For thermal units, variable cost information are provided, while fixed cost information were again obtained from literature (Kost et al.,, 2021; Schröder et al.,, 2013; Bundesnetzagentur,, 2022).
An additional challenge is the mapping of generators to corresponding network nodes. The grid data specify the broad category of generating units at each node (e.g., thermal, hydro, external injections) and their nominal capacities. While renewable resources could be mapped to the seller data by matching nominal capacities, the grid data lacks additional information regarding thermal plant types. We developed a heuristic to map thermal generators in the grid to their operational characteristics (such as minimum uptime constraints) (Schmitt,, 2023). Firstly, we match identifiers with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur to obtain each generator’s broad thermal generation type (e.g., lignite, gas, etc.). As nominal capacities do not match between grid and seller data, and seller data have a higher granularity (e.g., Gas CCGT, Gas OCGT, etc.), we solve an integer program to assign each entity in the grid model to exactly one corresponding seller of the same plant type, with the objective to minimize the aggregated differences in capacities and number of units per plant type. We provide this integer problem and the outcome of the generation matching in A While this assignment might not be exact, we argue that smaller mistakes in the assignment of generators to nodes have little impact on the more general questions raised in our paper that address the benefits of different zonal and nodal configurations. We provide our code files and a data description in an online appendix. The data are accessible through the ENTSO-E website.
3.2 Experimental Design
We implemented the models in Python 3.8 using Gurobi to solve optimization problems. We conducted experiments for eight weeks in 2009, as suggested by the ENTSO-E report (ENTSO-E,, 2022), for the German bidding zone. To maintain computational feasibility of computing hourly allocations and prices, we permitted a MIP gap of 5% for mixed-integer programs. Table 1 summarizes our experimental design and the scope of our analysis.
We first employ different allocation rules in accordance with the ENTSO-E report. These configurations range from a national single-zone model to a fully nodal model, as well as the ACER proposals for two to four bidding zones ACER, 2022b . Each allocation is associated with generation costs as well as redispatch costs to obtain a physically feasible outcome. After computing an allocation, different pricing rules, such as IP, CH, or Join pricing, can be applied. Each pricing rule implies different price levels as well as GLOCs, LLOCs, and MWPs for market participants. Prices were capped at EUR 100 per MWh to limit the impact of outliers. We also consider the Euphemia allocation and prices as the current implementation in European markets.
Calendar Weeks of 2009 Allocation Rules Pricing Rules Focus Variables
04 National IP Generation Costs
08 2 Zones (k-means) CH Redispatch Costs
11 2 Zones (spectral) Join Prices
15 3 Zones GLOCs
16 4 Zones LLOCs
21 Nodal MWPs
31 Euphemia
48
Table 1: Overview of experiments
4 Results
In this section, we summarize the main results of our analysis. We start with comparing average generation and total costs, before discussing price levels, GLOCs, LLOCs, and MWPs. We report aggregate statistics for all weeks under consideration and further pick the days of February 18, March 11, and November 23 for an in-depth analysis of representative days with low, medium, and high prices, respectively.
4.1 Generation and Redispatch Costs
Table 2 includes an overview of the average daily generation and redispatch costs (both for cost-minimizing and volume-minimizing redispatch) for the allocation rules under consideration.
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
Generation 33,507.00 33,507.68 33,507.37 33,516.13 33,510.40 36,008.27
Min-Cost Redispatch 5,124.61 5,120.69 5,127.95 5,117.03 5,145.39 0.00
Minimum Total Costs A 38,631.61 38,628.37 38,635.32 38,633.16 38,655.79 36,008.27
Min-Volume Redispatch 8,195.76 8,175.75 8,178.12 8,204.01 8,175.20 0.00
Minimum Total Costs B 41,702.76 41,683.43 41,685.49 41,720.14 41,685.60 36,008.27
Table 2: Average Daily Costs
The numbers reported in Table 2 have realistic orders of magnitude as compared to those provided by the German Bundesnetzagentur (Bundesnetzagentur,, 2024). However, our goal is not to come up with a precise prediction of costs for 2025, but to obtain a relative comparison between different zonal configurations and nodal pricing. Adding network constraints cannot lead to lower objective function values in the cost minimization. As the national allocation rule neglects all transmission constraints, their generation costs are a lower bound for all other allocation rules. On average, the suggested zonal configurations require only slightly more generation costs, suggesting that adding few transmission constraints to the computation has little impact on the cost. In contrast, the nodal allocation rule includes the entire transmission network. Obviously, higher generation costs are required to reach an allocation with feasible power flows. On average, generation costs increase by 7.46% after including all network constraints for the nodal model compared to the national configuration before redispatch. After redispatch, the national configuration with minimum cost redispatch has 7.29% higher costs than the nodal model, and 15.8% higher costs with minimum volume redispatch. Note that we assume that all generators can be redispatched in our model, such that redispatch costs in reality might be higher. In comparison, the average daily redispatch costs in Germany between 2020 and 2022 amount to roughly kEUR 7,200 (Bundesnetzagentur,, 2022). The average daily redispatch costs in Germany in 2022 were even at around kEUR 11,600, having increased by almost 100% compared to 2021. We assume zero redispatch for a nodal system that satisfies $\Psi^{DC}$, although in practice there will always be some minimal level of redispatch due to changing weather conditions and the simplified representation of network flows in the DCOPF.
4.1.1 Welfare losses with Euphemia - Impact of Non-Uniform Pricing
Euphemia maximizes welfare subject to linear and anonymous prices, which is different from unconstrained welfare maximization. In our analysis, the average daily welfare loss associated with Euphemia compared to the national configuration (without any network constraints) is 0.13% – or EUR 40,587.32 in absolute terms. This is comparable with previous estimates of 0.05% as a relative loss (Meeus et al.,, 2009), and with an absolute loss of EUR 130,000 for the entire SDAC region reported by All NEMO Committee, 2023a . Disregarding market coupling, the cuts introduced by the Euphemia algorithm to get to linear prices do not deteriorate the outcome by a lot. After redispatch, Euphemia suffers an average daily loss of 0.18% in total costs (or EUR 60,942.65) compared to the national allocation rule and a loss of 1.49% (or EUR 525,587.01) compared to the nodal allocation rule. We argue that the national, zonal, or Euphemia models differ only marginally in terms of allocation, and only a transition to a nodal allocation rule would substantially affect the average generation and redispatch costs.
4.2 Price Levels
Table 3 provides average prices over all hours of the weeks under consideration, while Table 4 illustrates median prices and Table 5 price standard deviations for all computed prices under the respective allocation and pricing rule.
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 42.68 42.54 42.04 42.20 43.30 44.34
CH 37.19 43.41 37.23 37.23 37.25 43.43
Join 40.85 42.55 42.12 42.29 41.44 43.48
Euphemia 43.41
Table 3: Average Prices
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 28.56 28.55 28.55 28.53 28.59 45.82
CH 30.81 30.81 30.81 30.81 30.81 46.52
Join 28.41 28.55 28.55 28.52 28.45 44.27
Euphemia 28.58
Table 4: Median Prices
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 19.85 19.71 19.40 19.62 20.05 19.44
CH 14.18 14.13 14.13 14.11 14.10 16.92
Join 21.53 20.21 19.96 20.17 21.78 21.08
Euphemia 20.63
Table 5: Price Standard Deviation
Note that the mean BZR price for the weeks under consideration was EUR/MWh 40.17 and the standard deviation EUR/MWh 18.51, computed bi-hourly based on a linear relaxation.
Average CH prices are slightly lower compared to the other pricing rules under consideration, with lower standard deviation and irrespective of the allocation rule. In contrast, median prices are at comparable levels across all pricing rules. Since CH prices are unaffected by pre-determined generator commitments, they tend to be lower and more stable. In particular, prices can be set by a cheaper – yet uncommitted – generator (Schiro et al.,, 2016). In contrast, average Euphemia prices tend to be slightly higher than zonal prices, while median prices are almost equal. This finding implies that the price-setting generator is usually identical for Euphemia and zonal prices, yet sometimes the Euphemia algorithm introduces a cut that disallows a generator that would otherwise be dispatched. As a result, a higher-price generator may set the price.
Price distributions are generally similar across the different zonal configurations for a specific pricing rule. Due to the simplified representation of the transmission network, cross-zonal flow constraints are rarely tight and thus zonal prices are identical to national (single-zone) prices in many hours. This observation suggests that splits into only a few zones, on average, have little impact on prices and thus provide few locational incentives. Table 6 illustrates that, on average, no major price discrepancies between zones can be expected. The zones are depicted in Figure 1.
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones
Zone 1 42.68 (19.85) 41.91 (19.47) 42.19 (19.64) 41.41 (19.05) 42.80 (20.31)
Zone 2 43.17 (19.93) 41.88 (19.16) 42.75 (19.81) 44.45 (20.66)
Zone 3 42.46 (19.95) 42.97 (19.60)
Zone 4 42.99 (19.56)
Table 6: IP Zonal Prices Average and Standard Deviation per Zone
(a) 2 Zones (k)
(b) 2 Zones(s)
(c) 3 Zones
(d) 4 Zones
Figure 1: Zone Configurations
A more in-depth examination of zonal pricing reveals the impact of cross-zonal flows in equalizing prices across different zones. Specifically, while southern zones face a relative shortage of generation capacity compared to demand, the capacity of cross-zonal transmission lines is typically high enough to transmit electricity from northern zones. Our sensitivity analysis revealed that if cross-zonal flows were limited or prohibited, zonal prices can display differences of more than 30 EUR/MWh between northern and southern regions on some days. The small differences in zonal prices observed in Table 6 are due to the cross-zonal transmission capacity (calculated as the sum of thermal limits of all cross-zonal lines) that equalizes prices across zones.
Under nodal pricing (ca. 40,000 prices per day for 1670 nodes – compared to, e.g., 24 national prices per day) more price outliers were observed than under zonal pricing. For instance, 0.08% of IP prices had to be capped at EUR 100 per MWh under the zonal configurations, compared to 2.26% under nodal pricing. Without capping, the maximum observed IP and Join price would be as high as EUR 70,000 per MWh for an hour of extreme scarcity. Note that SDAC also imposes a price cap of EUR 5,000 per MWh for such periods. CH prices, in contrast, are less prone to outliers, peaking at EUR 416 per MWh. Thus tight flow constraints can increase price levels in the short run. The zonal allocation rules oversimplify the transmission network to an extent that nodal prices can hardly be lower, even without congestion. In contrast, when congestion is present in parts of the grid, nodal prices will be higher, resulting in higher average and median prices. Upon further analysis, it was found that nodal prices have high sensitivity to transmission line susceptances, a phenomenon previously described in Bichler & Knörr, (2023).
Figure 2 illustrates nodal IP, CH, and Join prices, averaged over all hours and sorted in ascending order. The dashed horizontal and vertical lines mark the 5th and 95th percentiles for prices and nodes, respectively. Additionally, the figure shows the average national prices and the national average including a uniform EUR/MWh price adder for redispatch costs. This price adder is calculated as the total redispatch costs under min-volume redispatch (ca. EUR 459.0 million) divided by the total demand (ca. 83.3 TWh), and equals 5.51 EUR/MWh. Notably, with redispatch taken into account, higher average prices are expected at some of the nodes. For the majority of nodes, however, the change in average prices is minor or even negative under nodal pricing.
Figure 3 maps these average prices to their geographical locations. As much of the electricity supply, particularly wind energy, is located in Northern Germany, we observe that average nodal prices tend be higher in Southern Germany. Generally, this price gap is moderate and provides locational incentives for market participants. The choice of pricing rule does not seem to have a major effect on average price levels.
We observe the highest prices in Northern Central Germany in a region with a high share of expensive biomass energy and the lowest prices in the Alps region with inexpensive hydro energy. The associated nodes are relatively isolated with frequent congestion. This behavior might follow from imprecise estimations of generation costs and line limits described in Section 3, yet price outliers can also be observed in the ENTSO-E report (ENTSO-E,, 2022).
(a) IP Prices
(b) CH Prices
(c) Join Prices
Figure 2: Sorted Nodal Prices [EUR/MWh]
(a) IP Prices
(b) CH Prices
(c) Join Prices
Figure 3: Average Nodal Prices [EUR/MWh]
(a) IP Prices
(b) CH Prices
(c) Join Prices
Figure 4: Standard Deviation Nodal Prices [EUR/MWh]
Nodal prices typically vary across locations for each hour, and thereby set locational incentives for flexible units or storage. As seen in Figure 4, standard deviations of nodal prices resemble each other across nodes and pricing rules. Figure 5 illustrates the distribution of hourly standard deviations of prices across nodes, with mean and median standard deviations as solid and dashed lines, respectively. The nodal price variances can be substantial. For example, in the extreme first hour of 2009/11/24, IP prices vary across the country between EUR/MWh 3.45 and EUR/MWh 100. This sets incentives for demand response and short-term adjustments in particular in energy-intensive industrial production. Persistent high nodal prices signal transmission bottlenecks and direct investments in grid expansion.
(a) IP Prices
(b) CH Prices
(c) Join Prices
Figure 5: Histograms of Hourly Standard Deviations Across Nodes [EUR/MWh]
(a) IP Prices
(b) CH Prices
(c) Join Prices
Figure 6: Histograms of Nodal Standard Deviations Across Hours [EUR/MWh]
A primary goal in the design of new bidding zones is the reduction of price variance within a zone (ACER, 2022b, ), which signals congestion. Tables 5 and 6 indicate that the standard deviation of zonal prices does not decrease with the number of zones. However, varying prices can have two main reasons: congestion and temporal changes in supply and demand. Figure 5 considers hourly standard deviations of prices across all nodes, meaning it illustrates the price variance based on grid congestions in each hour, which we refer to as congestion-based standard deviation. In contrast, Figure 6 includes histograms of the standard deviation of nodal prices across all hours, meaning it illustrates the price variance based on temporal effects of supply and demand at each node, which we refer to as time-based standard deviation. Together, Figures 5 and 6 represent a composition of the standard deviations reported for nodal prices in Table 5. Note that prices at a single node can vary substantially over time due to changes in supply and demand. This time-based standard deviation exceeds the congestion-based standard deviation. For example, IP prices exhibit a mean congestion-based standard deviation of EUR/MWh 12.09, compared to a mean time-based standard deviation of EUR/MWh 16.95. These numbers provide evidence that temporal fluctuations of supply and demand have greater impact on price variance compared to congestion.
Tables 7 and 8 provide a more detailed variance decomposition for IP prices under all zonal configurations. In particular, Table 7 indicates the congestion-based standard deviation of nodal IP prices, averaged over all nodes assigned to a zone. Equivalently, Table 8 summarizes the time-based standard deviations grouped by zone. Specifically, the time-based standard deviation always exceeds the congestion-based standard deviation, implying that temporal effects drive price variance more than congestion effects. At the same time, the time-based standard deviation varies only little across allocation rules, suggesting that nation-wide fluctuations in supply and demand, on average, affect all price zones equally. In contrast, the congestion-based standard deviation has a slightly decreasing trend with more zones, indicating that zonal splits can reduce price variance at least for some of the zones (e.g., Zone 3). This effect, however, is not consistent across all zones, implying that the cross-zonal lines fail to encompass all congested lines. Consequently, the computed dispatch does not account for the remaining intra-zonal congestion, necessitating costly redispatch under all zonal models, as shown in Table 2. In a complementary paper, we demonstrate that the proposed zone configurations lack stability and fail to effectively segregate nodes along congested lines As a result, the congestion-based standard deviation exhibits minimal reduction as the number of zones increases, and redispatch costs remain high.
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones
Zone 1 12.09 12.29 12.31 11.51 12.52
Zone 2 10.45 10.53 12.47 10.97
Zone 3 6.42 6.44
Zone 4 12.85
Table 7: Average Congestion-Based Standard Deviation of Nodal IP Prices
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones
Zone 1 16.95 16.92 17.04 16.46 17.47
Zone 2 16.99 16.83 17.40 17.22
Zone 3 16.86 16.70
Zone 4 16.11
Table 8: Average Time-Based Standard Deviation of Nodal IP Prices
4.3 GLOCs, LLOCs, MWPs
The average daily GLOCs, LLOCs, and MWPs are summarized in Tables 911, respectively.
in EUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 2,743,625.97 2,715,067.90 2,581,180.46 2,665,434.36 3,073,576.90 6,667,208.89
CH 141,338.05 141,785.33 142,372.81 153,430.35 145,411.76 361,031.42
Join 2,784,666.32 2,773,482.10 2,680,900.33 2,719,728.69 3,115,596.04 4,665,204.10
Euphemia 2,772,814.13
Table 9: Average Daily GLOCs
in EUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 0.00 0.00 0.00 0.00 0.00 0.00
CH 33,683.28 32,956.96 33,554.87 40,266.34 35,536.04 179,887.45
Join 989.05 812.56 1,547.75 1,795.06 1,057.30 41,476.85
Euphemia 0.00
Table 10: Average Daily LLOCs
in EUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 19,812.81 25,662.91 21,626.78 28,348.30 21,558.54 338,926.64
CH 37,345.65 37,658.44 37,975.21 45,054.94 39,156.73 102,818.68
Join 766.2 1,308.67 1,446.57 1,094.99 859.76 23,840.19
Euphemia 0.00
Table 11: Average Daily MWPs
By definition, CH prices minimize GLOCs, and on average, they are substantially lower compared to both IP and Join prices, irrespective of the chosen allocation rule. Similar orders of magnitude for GLOCs are observed across all zonal allocation rules, including Euphemia. In contrast, nodal pricing results in the highest GLOCs, aligning with the higher average prices presented in Table 3. Higher prices imply greater forgone profits for generators and increase GLOCs. Generators in nodal markets face stronger incentives to deviate from the optimal outcome, necessitating penalties – a common practice in many U.S. markets (Bichler et al.,, 2022).
As a natural subset of GLOCs, LLOCs follow a similar pattern, with the highest values occurring under nodal pricing, once again correlated with higher average prices. IP pricing ensures zero LLOCs, as discussed in Section 2.3.2, ensuring reliable congestion signals where price differences between nodes only arise during network congestion. In this case, the price difference precisely reflects the marginal value of transmission capacity. In contrast, CH pricing consistently results in the highest LLOCs, supporting observations that congestion signals may be flawed (Schiro et al.,, 2016). This is especially relevant under a nodal allocation rule considering all transmission lines. The Join pricing rule yields substantially fewer LLOCs than CH pricing, indicating a superior quality of congestion signals.
In terms of average MWPs to compensate generator losses, only the Euphemia algorithm generates zero MWPs, albeit at the expense of welfare losses. An interesting observation is that the welfare losses of Euphemia – as discussed in Section 4.1 – exceed, on average, the MWPs required under a national welfare-maximizing allocation, irrespective of the pricing rule. This implies that the additional generated welfare could potentially cover the MWPs under a national allocation, resulting in a net welfare gain. Under nodal pricing, MWPs tend to increase compared to zonal pricing, particularly for IP pricing. This can be attributed to the fact that under a national allocation rule, there exists a single hourly IP price set at the variable cost of the marginal generator. With positive fixed costs and constant variable costs, such a generator will inevitably incur a loss and necessitate MWPs under IP pricing. In contrast, nodal pricing involves several hourly IP prices and price-setting generators across the network, leading to higher overall MWPs. Conversely, the Join pricing rule generates lower MWPs than IP and CH pricing, constituting a negligible share of the total costs outlined in Table 2.
In summary, GLOCs, LLOCs, and MWPs follow similar patterns across different zonal configurations, but increase with nodal pricing. Concerning pricing rules, IP, CH, and Euphemia each minimize a specific class of lost opportunity costs. However, the Join pricing rule stands out by striking a remarkable balance between low MWPs and LLOCs. At the expense of slightly increased GLOCs, Join prices facilitate marginal side-payments and maintain a high quality of congestion signals.
4.4 Analysis of Representative Days
To corroborate our findings, we examine three selected days in greater detail. Specifically, we analyze November 23 for its low price levels, March 11 for medium price levels, and February 18 for high price levels.
The generation costs for each of these days are presented in Table 12. On 2009/11/23, characterized by warm and windy conditions with low demand, generation costs were minimal, and no congestion occurred in the zonal models, resulting in identical allocations. In contrast, February 18, as a cold and cloudy day, saw substantially higher generation costs, but cross-zonal flow constraints had little impact on the allocation. Across all days, the shift from zonal to nodal pricing led to a similar increase in generation costs.
Table 13 indicates that minimum cost redispatch costs are not correlated with the generation costs presented in Table 12. This observation underscores that zonal models lack a representation of transmission bottlenecks, necessitating redispatch regardless of the zonal configuration or day.
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 26,672.26 26,672.26 26,672.26 26,672.26 26,672.26 29,593.76
2009/03/11 30,494.25 30,507.06 30,502.60 30,507.54 30,496.97 33,068.16
2009/02/18 45,481.38 45,483.18 45,483.04 45,666.68 45,486.15 48,718.75
Table 12: Generation Costs: Nov 25, Mar 11, Feb 18
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 3,457.50 3,457.50 3,457.50 3,457.50 3,457.50 0.00
2009/03/11 2,969.30 2,989.12 2,955.14 2,965.60 2,990.37 0.00
2009/02/18 3,649.39 3,664.49 3,639.52 3,561.45 3,596.63 0.00
Table 13: Min-Cost Redispatch Costs: Nov 25, Mar 11, Feb 18
In summary, as outlined in Table 14, the nodal allocation rule consistently yields lower total costs than any zonal configurations. Despite modest cost savings, nodal allocation enables locational price signals and incentivizes a long-run equilibrium.
On 2009/11/23 and 2009/03/11, the allocation obtained from the Euphemia algorithm is identical to the national configuration, and no welfare loss occurs. On 2009/02/18, costs under the Euphemia algorithm are slightly higher by 0.36%. This validates our assertion that differences between zonal models – including Euphemia – are minor and nuanced, while nodal pricing constitutes a more substantial change in generation and redispatch costs.
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 30,129.75 30,129.75 30,129.75 30,129.75 30,129.75 29,593.76
2009/03/11 33,463.55 33,496.18 33,457.73 33,473.14 33,487.34 33,068.16
2009/02/18 49,130.77 49,147.67 49,122.56 49,228.13 49,082.78 48,718.75
Table 14: Total Costs: Nov 25, Mar 11, Feb 18
Figure 7 illustrates the locational IP, CH, and Join prices for the three selected days, averaged over 24 hours. As previously discussed, the prices on these days exhibit a positive correlation with generation costs, with the exception of Northern Central and Southern Germany with consistent price outliers due to local supply structures. Notably, even though at lower price levels, a more distinct north-south price gradient can be observed on 2009/11/23 compared to 2009/03/11. Specifically, on 2009/11/23, demand is low nationwide, yet inexpensive electricity from the north cannot be transmitted to the south, leading to uneven price levels. In contrast, on 2009/03/11, with higher demand but a balanced grid, prices are uniform across the country. This phenomenon holds for all three pricing rules. The simplified zonal models fail to detect any transmission bottlenecks on 2009/11/23, i.e., cross-zonal constraints are not tight and there are no price variations between zones. Consequently, if the zone configurations are suboptimal, zonal prices do not signal scarcity appropriately. B includes an analysis of hourly prices observed on the three days.
(a) 2009/11/23 - Avg IP Prices
(b) 2009/11/23 - Avg CH Prices
(c) 2009/11/23 - Avg Join Prices
(d) 2009/03/11 - Avg IP Prices
(e) 2009/03/11 - Avg CH Prices
(f) 2009/03/11 - Avg Join Prices
(g) 2009/02/18 - Avg IP Prices
(h) 2009/02/18 - Avg CH Prices
(i) 2009/02/18 - Avg Join Prices
Figure 7: Average Nodal Prices
Table 8 summarizes the GLOCs, LLOCs, and MWPs for each selected day. On the low-demand day of 2009/11/23, a Walrasian equilibrium as in Definiton 1 is attainable for the zonal models, resulting in zero GLOCs, LLOC, and MWPs. Crucially, all three pricing rules successfully obtain these equilibrium prices. The Walrasian equilibrium, however, can only persist before redispatch is considered and ceases to exist when nodal transmission constraints are taken into account.
GLOCs, LLOCs, and MWPs do not necessarily increase with average price levels. For nodal pricing, they remain relatively consistent across all days. There is a substantial increase for zonal pricing on the medium-price day of 2009/03/11, compared to the high-price day of 2009/02/18. This indicates that incentives to deviate hinge more on the quality of the price signal than the actual price level.
The simplification inherent in zonal models, which leads to redispatch, does not necessarily imply better economic properties of prices. For example, on 2009/03/11 and 2009/02/18, zonal IP and Join prices result in higher GLOCs than nodal pricing. Specifically, a few zonal prices are insufficient to incentivize many market participants to adhere to the allocation.
As discussed above, CH prices always minimize GLOCs, and IP prices always yield zero LLOCs. Although GLOCs are reduced by CH pricing compared to both IP and Join pricing, their LLOCs, and often MWPs, experience a slight increase. In practical terms, this implies that congestion signals are more distorted and higher side-payments are required. In contrast, the Join pricing rule achieves a favorable trade-off between LLOCs and MWPs, requiring only slightly higher penalties (corresponding to GLOCs) than IP pricing.
in kEUR IP CH Join
National GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
2 Zones (k) GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
2 Zones (s) GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
3 Zones GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
4 Zones GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
Nodal GLOC 1,277.13 272.47 1,492.66
LLOC 0.00 147.44 26.79
MWP 167.32 94.76 34.70
(a) GLOC, LLOC, MWP for 2009/11/23
in kEUR IP CH Join
National GLOC 4,323.93 43.73 4,323.93
LLOC 0.00 10.68 0.00
MWP 0.00 9.04 0.00
2 Zones (k) GLOC 8,009.06 56.54 8,009.06
LLOC 0.00 23.49 0.00
MWP 0.00 23.03 0.00
2 Zones (s) GLOC 5,419.69 52.08 5,419.69
LLOC 0.00 20.17 0.00
MWP 0.00 17.05 0.00
3 Zones GLOC 7,097.32 57.03 7,097.32
LLOC 0.00 23.98 0.00
MWP 0.00 23.53 0.00
4 Zones GLOC 5,152.01 46.45 5,152.01
LLOC 0.00 13.35 0.00
MWP 0.00 11.10 0.00
Nodal GLOC 1,146.96 178.49 1,159.88
LLOC 0.00 96.05 37.93
MWP 94.90 29.73 22.33
(b) GLOC, LLOC, MWP for 2009/03/11
in kEUR IP CH Join
National GLOC 1,157.36 290.42 1,177.70
LLOC 0.00 30.30 9.38
MWP 59.64 22.53 9.29
2 Zones (k) GLOC 1,798.28 288.22 1,772.59
LLOC 0.00 28.16 9.58
MWP 132.61 18.85 9.50
2 Zones (s) GLOC 1,430.45 285.19 1,455.50
LLOC 0.00 25.13 9.36
MWP 58.69 13.73 9.27
3 Zones GLOC 2,387.52 471.72 2,493.28
LLOC 0.00 170.36 9.20
MWP 27.32 164.81 9.02
4 Zones GLOC 1,778.31 291.19 1,860.59
LLOC 0.00 30.96 9.41
MWP 61.81 20.96 9.33
Nodal GLOC 1,024.32 274.54 1,495.84
LLOC 0.00 86.48 14.44
MWP 72.94 34.68 9.78
(c) GLOC, LLOC, MWP for 2009/02/18
Figure 8: Daily GLOCs, LLOCs, MWPs
4.5 Discussion
Our numerical experiments provide insights into different market clearing models, pricing rules, and their economic implications on day-ahead markets. Note that the final dispatch is affected by additional factors such as intraday trading or long-term capacity allocations. Nodal pricing tends to result in slightly higher average prices compared to zonal pricing, irrespective of the chosen zonal configuration. However, Transmission constraints can be violated and necessitate costly redispatch measures. As a result, a zonal pricing will always pick an inefficient dispatch, and even if redispatch is conducted at minimal costs, the total system costs will be substantially higher than under nodal pricing. Given the steep increase of redispatch costs in recent years, this effect is likely to magnify in the future.
Interestingly, the choice of zone configuration did not substantially impact prices. All zonal configurations suggested in the Bidding Zone Review simplify the network to an extent that opting for two, three, or four zones yields similar results regarding the average prices and the price standard deviation. In contrast to nodal pricing, zonal pricing leads to substantial redispatch and higher total costs. While nodal prices may vary between nodes, these differences incentivize short-term demand response on nodes where this is most helpful for overall system stability. If such price differences persist over more extended time periods, they also set investment incentives for generators and grid expansion.
Given the absence of Walrasian equilibria, any pricing rule – regardless of zonal or nodal pricing – must balance economic properties such as welfare gains and the participants’ lost opportunity costs. The currently used Euphemia algorithm leads to welfare losses and is computationally expensive but avoids paradoxically accepted bids. Our experiments suggest that non-uniform pricing rules achieve low MWPs for paradoxically accepted bids while maintaining the optimal allocation and scaling in polynomial time. Our findings thus provide support for non-uniform pricing as they are currently being discussed in the context of the Capacity Allocation and Congestion Management (CACM) Regulations (All NEMO Committee, 2023b, ).
CH pricing minimizes GLOCs, but our results confirm previous findings that congestion signals are distorted (evidenced by high LLOCs) and high MWPs are necessary. IP pricing, the prevailing rule in U.S. markets, ensures effective congestion signals but requires very high MWPs. The Join pricing rule strikes a balance between LLOCs and MWPs, ensuring low side-payments and good congestion signals. GLOCs are less of a concern, because already today, in U.S. markets with IP pricing, such incentives to deviate from the computed dispatch are mitigated by penalties for deviations.
5 Conclusion
The discussion surrounding zonal and nodal pricing has a long history. The European electricity market has employed zonal pricing for many years, often with large, nationwide price zones. However, the ongoing energy transition requires a reevaluation of price zones. Based on recent data released by ENTSO-E in the context of the EU Bidding Zone Review process, we compare various allocation and pricing rules for the German power market. Our findings indicate that, in terms of system costs and price signals, a nodal allocation rule with non-uniform pricing leads to the lowest total costs. Costly redispatch is mitigated, and the average prices increase only slightly. Additionally, pricing rules such as Join pricing require very low side-payments, while efficiently signaling transmission bottlenecks. In contrast, the proposed zonal configurations only differ marginally regarding system costs and the differences in average prices across zones are low. Interestingly, the effect of zonal splits on price standard deviations is low, and we do not find evidence for significant reductions in redispatch costs resulting from such splits, all other factors held constant.
Acknowledgements
We would like to thank Karsten Neuhoff and Maximilian Rinck for many helfpul comments. This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. BI 1057/10.